If tanx = 3/7 find the exact value of cosx.
Answer by lwsshak3(11628)

If tanx = 3/7 find the exact value of cosx.
assume x=acute angle
hypotenuse of reference right triangle=√(3^2+7^2)=√(9+49)=√58
tanx=sinx/cosx=3/7
cosx=7/√58=7√58/58